The general form of the equation of a straight line can be expressed as Ax + By + C = 0 with a condition that A and
B cannot both be zero. This equation can also be expressed as: y=
−(
1
Expert's answer
2019-09-18T07:36:01-0400

As y=mx + b which is Slope-Intercept Form of equation of a straight line.

For example let's convert 6x - 2y - 3 = 0;

We are heading for y=mx + b;

Start: 6x - 2y - 3 = 0;

Move all except y to the right: -2y = -6x + 3

Divide all by (-2): y = 2x - 5/2

And done!

(Note that m = 2 and b = -5/2 OR m = A/B and b = C/B)